# Jon Danielsson London School of Economics Andreas Uthemann

### v4

@@ 4.2 The risk of authorities losing control
As the public and private sectors increase their use of AI, the authorities must carefully consider how they adopt AI. They might conclude that AI should only be used for advice, not decisions, with human beings in the loop to avoid undesirable outcomes. However, that might not be as big a distinction as the authority thinks. The AI engine will have its internal representation of the financial system, where its understanding might not be intelligible to its human operators. When we then use that AI to scan the system for vulnerabilities and run scenarios to evaluate the impact of alternative regulations or courses of action for crisis resolution, we might have no choice but to accept its advice, especially when presented as a choice between something that appears sensible and a potentially disastrous alternative. When optimising, the AI may even choose to present alternatives in that manner so as not to risk having the operator make inferior choices.
4.3 Trust
While an authority might not wish to get to that point, its use of AI might end up there regardless. As we come to trust AI analysis and decisions and see how well it performs in increasingly complex and essential tasks, it may end up by stealth, where the authorities do not want to be. That arises because of how AI creates trust. There are many examples where technology was initially met with scepticism. As it performs, technology is then increasingly accepted, as we have seen with AI landing passenger aeroplanes, AI-guided surgery and self-driving cars. Even if we only allow AI to execute the simplest tasks inside the financial authorities, as it starts performing at the same level or better than its human counterparts, but at a much lower cost, we can expect AI to become increasingly welcomed. In other words, its very success creates trust. As that trust is earned on relatively simple and safe repetitive tasks and seeing AI succeed in more complex jobs, it will be asked to take on ever more sophisticated tasks.
@@ 4.3 Trust
As trust builds up, the critical risk is that we become so dependent on AI that the authorities cannot exercise control without it. In a sense, AI optimises to become irreplaceable. By then, turning the AI engine off may be impossible because it performs vital functions, and the risk of disastrous outcomes might be deemed unacceptably high. Eventually, we risk becoming dependent on a system for critical analysis and decisions we don’t entirely, or even partially, understand.
4.4 Six criteria for AI use in financial policy
@@
The conceptual challenges in Section 2 and the various issues surrounding AI take us to 6 criteria for evaluating AI use in financial policy.
@@
Can AI be given clear objectives and monitor its actions in light of those objectives, or are they unclear?
@@
Would a human functionary have the authority to make decisions, does it require committee approval, or is a fully distributed decision making process brought to bear on a problem?
@@
5 The channels for how AI can destabilise the financial system
While we anticipate that AI will be broadly beneficial to the financial system, it may also increase systemic financial risk, defined as a serious financial crisis that has a high chance of causing an economic recession. The costs of systemic crises are very high, in the trillions of dollars for large economies, as noted by Barnichon et al. (2022). As such crises are usually caused by excessive low-quality lending, amplified by high leverage and asymmetric information, they should be easily preventable as macroprudential regulations aim to do. It is not so in reality. The almost infinite complexity of the financial system makes it very hard to design effective regulations that simultaneously prevent excesses and do not hamper economic growth too much. On top of that, because the economy is usually growing rapidly before a crisis, there is strong political opposition to any measures that might threaten that growth.
The reason why AI can increase systemic risk is, perhaps paradoxically, its very strength and efficiency. In the words of Bengio et al. (2023), “Compared to humans, AI systems can act faster, absorb more knowledge, and communicate at a far higher bandwidth. Additionally, they can be scaled to use immense computational resources and can be replicated by the millions.” The financial system has a large number of potentially destabilising feedback mechanisms that are created by the interaction of actors and the significant resources that are trying to maximise profits in regular time and survival during stress. The combination of AI that excels at discovering optimal strategies, the system’s complexity, highly resourced actors and the inherent destabilising factors already prevalent in the financial system viciously reinforce each other.
We build on extant work on AI safety, like Weidinger et al. (2022), Bengio et al. (2023) and Shevlane et al. (2023), which identifies several societal risks arising from AI use, including malicious use, misinformation and loss of human control. When augmenting that, we propose four channels for how AI use may destabilise the financial system, all of which follow from the AI issues discussed in Section 3 and the conceptual challenges in Section 2.
5.1 Malicious use of AI
The first channel arises from how the operators of financial AI use them to optimise against the system for malicious purposes. Financial institutions and their staff deploy considerable resources in their pursuit of profit and usually are not very concerned about the wider social impact of their activities. They can change the system in ways that benefit them while not being detectable by others, as we have seen many times in the past. AI provides such actors with new opportunities, either via adversarial attacks by feeding particular data into its training algorithms or influencing network design. That might not even be necessary since it is straightforward to exploit for private, even illicit, gain gaps in how AI sees its responsibilities.
One way such malicious use might occur is by manipulating governance processes. The financial authorities and internal control aim to align the interests of system participants to the organisation that employs them and society. As AI is particularly effective in finding profitable loopholes and amplifying vulnerabilities, it can facilitate misbehaviour that, while legal, is damaging both to society and even the institution employing the AI.
The operator of the AI could take that one step further and use it to deliberately create market stress, which, of course, is profitable to those forewarned. We have seen many examples of such conduct in the past, and a key purpose of securities and macroprudential regulations is to prevent such behaviour.
And finally, we have those aiming to manipulate the system for illegal purposes, like rogue traders, criminals, terrorists and nation-states. The inherent vulnerability of the financial system to adversarial attacks and the ease of manipulating AI engines create potential for particular attack vectors that can be impossible to identify ex-ante. One example comes from Hubinger et al. (2024) who identify what they call “sleeper agent” that usually acts as expected but gives deceptive answers when given special instructions. Here, AI engine corruption becomes another form of cyber attack. With such an attack vector in place, an AI engine becomes a tool for well-resourced hostile agents that might be difficult or impossible to prevent in real time, facilitating nation-state attacks on critical infrastructure. That is particularly relevant when a superior ability to identify optimal timing strategies allows them to solve the problem of double coincidence, as discussed in Daníelsson et al. (2016), creating heightened system fragility, perhaps by manufacturing a liquidity crisis as an attack amplifier.
5.2 Misinformed use and overreliance on AI
The second channel emerges from humans using AI incorrectly. That is particularly relevant when AI is used in domains with unclear objectives, which is very common in the financial system. As events become less frequent and more serious, data-driven decisions increasingly involve extrapolation, interpretation, intuition and nuance, none of which AI in its current form is good at. As we discuss on Page 2.4, many crises have been resolved by very creative solutions that had never been used before.
Because successful AI architectures are data-driven engines, their analysis is predominantly founded on statistical regularities. Such AI engines are designed to provide advice and responses to prompts, even if they have very low confidence about the accuracy of the answer. They can even make up facts or present arguments that sound plausible but would be considered flawed by an expert, both instances of the broader phenomenon of AI hallucination. The risk is that the AI engines will present confident recommendations about outcomes they know little about. To overcome that, the engines should be required to provide an assessment of the statistical accuracy of their recommendations. Here, it will be helpful if the authorities overcome their frequent reluctance to adopt consistent quantitative frameworks for measuring and reporting on the statistical accuracy of their data-based inputs and outputs.
5.3 AI misalignment and control avoidance
The problem of misalignment is particularly difficult when several AI engines interact. For example, market manipulation is most successful when several investors collude. Similarly, the payoff of many trading activities, such as carry trades and short sales, increases with the number of investors participating in them. A particularly damaging consequence of the strategic complementarities prevalent in the financial system relates to the objectives of financial institutions. They maximise profit most of the time, perhaps 999 days out of a thousand, but optimise for survival during stress. The behaviour on that one day out of a thousand is particularly damaging when financial institutions seek safety by withdrawing liquidity, which can lead to destabilising dynamics such as fire sales, bank runs, and credit crunches. The very high level of AI performance can, perhaps paradoxically, increase the likelihood of damaging coordinated behaviour. They might find better strategies for exploiting strategic complementarities for private gain, at the risk of lowering market quality or even threatening financial stability.
An example of AI collusion is Calvano et al. (2020), who find that independent reinforcement learning algorithms instructed to maximise profits quickly converge on collusive pricing strategies that sustain anti-competitive outcomes. It is much easier for AI to behave in this collusive way than humans, as such behaviour is very complex and often illegal. AI is much better at handling complexity and is unaware of the legal nuances unless explicitly taught or instructed.
The most difficult problems of misalignment arise due to the interaction of the various AI working throughout the industry. Long before computers got involved with the financial system, humans coordinated damaging behaviour due to the misalignment of individual short-term incentives with the long-term incentives of the same market participants and society at large. The crises of 1763, 1914 and 1929 (see, e.g. Daníelsson, 2022) are just two of many such examples, as is the tulip mania in the Netherlands in the 1630s, see Aliber and Kindleberger (2015). The use of computers amplifies the potential for such undesirable outcomes as two recent examples illustrate. The largest stock market crash in history happened in 1987 when global markets fell by 23% in one day. The reason was trading algorithms that, in a way not foreseen, coordinated in creating vicious feedback between prices collapsing and algorithm-induced selling, as documented by Gennotte and Leland (1990). Similarly, the quant crisis in the summer of 2007, as noted by Khandani and Lo (2007), was due to trading algorithms trained to sell on falling prices but not being told of the potential for aggregate feedback caused by trading algorithms in a number of funds coordinating on selling. Such factors have further been at the root of the various flash crashes. AI will likely amplify such problems of misalignment.
It is tough for the authorities to patrol a nearly infinitely complex system. They have limited resources and can only look at specific system parts. Once they are focused on particular behaviour, that risk might be contained, but what tends to happen is that the forces of instability emerge elsewhere. Risk is a bit like a balloon; squeeze it in one area, and it expands elsewhere. Indeed, it is almost axiomatic that instability arises where the authorities are not looking. And, in an almost infinitely complex system, there is plenty of scope for instability.
AI will help the authorities keep the system stable and aid the forces of instability. The question is, which of these two dominate? Instability, as it needs only to find one weakness, while the authorities must monitor the entire system. AI will be particularly good at identifying the weak points.
Meanwhile, the supervisor must not only identify the weak points but also monitor how financial institutions interact with each weakness in real time. A much more difficult computational task, and one where the more complex the system becomes, the harder the task is. Furthermore, the supervisors have fewer resources than the private sector. The regulator might even have to rely on private sector resources (compute, fitted models, data). The resulting power asymmetry may be impossibly hard to solve.
That means AI, either autonomously or at human instruction, would likely be particularly good at finding strategies for avoiding both internal controls in a financial institution and regulatory oversight, making the two-sided principal-agent problem we discuss in Section 2.5 particularly challenging. In other words, the more sophisticated AI become, the more difficult it will be to exercise authority over them. After all, the controllers need to understand the AI objectives to design the correct incentives. And given the current state of interpretability of foundation models, this is very challenging.
5.4 Resource concentration
Monoculture is an important driver of booms and busts in the financial system. As financial institutions come to see and react to the world in increasingly similar ways, the more they coordinate in buying and selling, leading to bubbles and crashes or, even worse, credit booms and credit crunches. Banks’ capital calculation is a particular driver of procyclicality, not the least via risk weights. Such procyclical behaviour has always been a feature of financial markets, long before the use of computer technology. Still, it is amplified by technology.
Finally, financial data vendors have concentrated considerably over the past few years, with only a handful left, such as S&P Global, Bloomberg and LSEG. They have unified databases with standardised labelling and graph structures linking disparate categories of data, including proprietary and not publicly available data, facilitating training. It is a concern that neither the competition nor the financial authorities appear to have fully appreciated the potential for increased systemic risk due to oligopolistic AI technology in the recent wave of data vendor mergers.
Taken together, AI-driven financial analytics is characterised by increasing returns to scale technologies with high entry costs, making an oligopolistic market structure highly likely. In such an environment, outsourcing analytics to one of those vendors might be the best, maybe even the only feasible solution for all but the largest financial companies. Anecdotal private sector evidence indicates that many financial institutions already prefer to outsource data-driven applications to specialists. A particular example is risk management as a service, RMaaS, with a leading vendor being BlackRock’s AI-fuelled Aladdin. The risk is harmonised beliefs and actions, driving procyclical behaviour and exposing users to the same blind spots.
6 Regulating AI
The financial system is one of the most regulated parts of the economy because of the strong externalities arising from financial activities. A particular challenge for macropru is that keeping individual parts or institutions of the system safe is not equivalent to keeping the system as a whole safe. In fact, these two objectives often come into conflict, and regulation needs to take a system perspective with all the complexity that ensues.
6.1 Regulation of private AI
As the private sector increasingly adopts AI, the supervisors will have to address both how AI changes the behaviour of the private sector institutions and also how to regulate the AI engines in use. Several issues emerge.
Both the public and private sectors will likely end up outsourcing analytics to the same handful of very large AI vendors. This blurs the regulator/regulated divide since if the private sector is acting on the same advice as the public sector, the nature of supervision will be different than today. The reason is that a vendor’s engine will likely have a single representation of the financial system, that is, of the stochastic processes that drive market outcomes. This means that advice provided to all users of the engine, in the private and public sectors, will share the same underlying view of the system. That has two consequences. The first is that it will harmonise beliefs, which is pro-cyclical. Furthermore, since the engine is just a model that maps a near infinitely complex system to a simplified version of it, if a private firm unknowingly engages in undesirable actions at the recommendation of the engine, the authority using the same engine might not identify that misconduct. To guard against both issues, the authorities should pay particular attention to problems of model diversity and robustness, aim to use different engines than regulated entities and, seek to have several vendors of AI analytics and take advice from several vendors on important questions of system stability.
While regulations can prevent that, it would then only come at the expense of lower performance.
Increased use of private sector AI can result in a new layer of deniability. Assigning legal responsibility for misconduct in the financial sector is already tricky. It will be harder when AI is used to make decisions. Suppose a human operator deliberately instructs AI to break the law for criminal or terrorist purposes or just turns a blind eye to the AI doing so as a byproduct of maximising profits. Even if detected, it might be easily explainable as an unintended and unexpected innocent behaviour. For example, while a rogue trader can be prosecuted, we cannot do that with a rogue AI trader. The institution employing such AI might profit significantly. Consequently, the increased use of AI in the private sector facilitates the job of those seeking to utilise AI for nefarious purposes by providing them with yet another level of denial, at least until the law and regulations catch up.
As the authorities ease into regulating a financial system that uses AI, a helpful approach might be to evaluate the safety of private AI against defined “benchmark tasks” that are well-understood and clearly defined everyday regulatory activities. While that might give rise to the private sector optimising against those benchmark tasks, constructive ambiguity in their design would likely significantly hamper such efforts.
Bengio et al. (2023) and Shevlane et al. (2023) in their analysis of the use of AI, propose avenues for benchmarking, such as checking for engine consistency, that is, whether standard inputs yield reasonable outputs, and that similar inputs yield equally similar outputs. The authorities could then execute regular scans for engines having unexpected capabilities and test them to see if they can engage in manipulative behaviour and can be transparent, via explainability, in how they reach conclusions. Ultimately, transparency, for example, by publishing sufficient details on architecture and training or even using open-source approaches, will allow outside scrutiny.
6.2 Regulation of public AI
The financial authorities act as agents for society and have to be seen as accountable to society. We have several safeguards built to achieve that, including judicial and parliamentary scrutiny. We have governance structures in place. As the authorities increasingly start using AI, new areas of accountability are raised.
We have several ways today to provide that accountability. The human supervisor who makes mistakes can be held accountable, trained, dismissed, or even prosecuted, as we have seen many times in the past. If AI is tasked with making the same type of supervisory decisions, how does one attribute accountability? While the supervisory structure has many individuals, where perhaps only one makes a mistake, the authority will use only one AI engine to make all supervisory decisions. That means mistakes become systemic, not idiosyncratic. However, the engine can’t be turned off if it is already in use, and retraining it won’t be easy. The authority will need to have a response plan in place for such eventualities.
When a human supervisor makes a mistake, or the regulated thinks a human supervisor made a mistake, the regulated can appeal and challenge the ruling in a court of law. That means a human supervisor will be called on to explain the logic behind a decision, and the authority will explain the guidelines given to the human supervisor. That will be more difficult when AI makes decisions. There will be no separation between individual decisions and the guidelines, and the regulatory AI may not be able to explain its reasoning or why it thinks it complies with laws and regulations.
As an engine becomes embedded, we get to the point where we can not turn the engine off because that might lead to an entire essential function in the supervisory apparatus left unaddressed. Ultimately, this means that the internal supervision of AI use in the regulatory agencies and its interaction with the legal system will require different policies than those used for current human supervisors.
A useful way to evaluate AI use in the public sector would be to test them for safety and performance on defined benchmark tasks, particularly activities the authority engages in. It could be asked to advise and rule on macro prudential actions and suggest how to respond to different hypothetical crisis scenarios.
@@ 7 Conclusion
The literature on AI has identified several areas where it can threaten society, and we augment those with issues arising in the financial system. We find five channels: malicious use of AI, misinformed use and overreliance on AI, AI misalignment, AI as an increasing return to scale technology, and AI evading control.
Table 1: Particular regulatory tasks and AI consequences
Task Data Mutability Objectives Authority Responsibility Consequences Fraud/Compliance Consumer protection Ample Very low Clear Single Mostly clear Small Microprudential risk management Routine forecasting Ample Very low Mostly clear Single Clear Moderate Criminality Terrorism Limited Very low Mostly clear Multiple Moderate Moderate Nation state atttacks Limited Full Complex Multiple & international Moderate Very severe Resolution of small bank failure Limited Partial Clear Mostly single Mostly clear Moderate Resolution of large bank failure Severe market turmoil Rare Full Complex Multiple Often unclear Severe Global systemic crises Very rare or not available Full Complex & conflicting Multiple & international Unclear even ex-post Very severe
@@ Appendix A Machine Learning (ML)
There are many excellent sources on the technical details of ML, see for example Hastie et al. (2009) and Murphy (2023). The ML algorithms that underpin the current generation of successful AI engines are based on neural networks and involve inferring a complex mapping from some input data to an output. [^10] The algorithms look for a function $f$ that “best” maps input $x$ to output $y$, where $f$ involve multiple nested layers, transforming the input data to extracting complex patterns. The techniques used to fit or train these types of models are referred to as deep learning. See Prince (2023) for an up-to-date overview of deep learning techniques. For example, for large language models (LLMs), the input is a sequence of what is called tokens, perhaps words, images or sounds, while the output is the next token in the sequence, such as a word in a sentence. [^11] Currently, the most successful architecture for foundation models is the transformer (Vaswani et al., 2023) used in OpenAI’s GPT models, Google’s PaLM and Meta’s LLaMA. While the transformer architecture dominates the current AI landscape, some alternative architectures for foundation models have recently emerged, e.g., state space sequence models (SSMs) (Gu et al., 2022) that are computational efficient and can model long-term dependencies in time-series data.

### v5

@@ 5.2 The authorities’ challenges
The leadership might conclude that AI should only be used for basic advice, not decisions, ensuring that humans are always kept in the loop to avoid undesirable outcomes. That might be wishful thinking and not as big a distinction as one might think. The AI engine will have its internal representation of the financial system, where its understanding might not be intelligible to its human operators. When we then use that AI to scan the system for vulnerabilities and run scenarios to evaluate the impact of alternative regulations or courses of action for crisis resolution, we might have no choice but to accept its advice, especially when presented as a choice between something that appears sensible and disastrous. When optimising, the AI may even choose to present alternatives in that manner so as not to risk having the operator make inferior choices, just like human advisors are prone to do.
@@
As trust builds up (Section 4.2), the critical risk is that we become so dependent on AI that the authorities cannot exercise control without it. In a sense, AI optimises to become irreplaceable. By then, turning the AI engine off may be impossible because it performs vital functions, and the risk of disastrous outcomes might be deemed unacceptably high. That is unlike today when we react to flash crashes with market stops. Once technology is much more prevalent and no longer localised, we may lose access to such simple tools. Eventually, we risk becoming dependent on a system for critical analysis and decisions we don’t entirely, or even partially, understand. Ultimately, this means that the internal supervision of AI use in regulatory agencies and its interaction with the legal system will require policies different from those used by current human supervisors.
5.3 Regulation of private AI
The authorities already regulate how the financial industry uses technology and models. Trading algorithms have to be validated and risk models approved, while the design of data centres, cloud use and outsourcing are very much the purview of the authorities. That implies models based on AI will similarly be regulated. However, because of the particular way they learn from the environment, which is different from more traditional statistical and ML techniques, the regulatory methodology will have to change.
One approach might be to evaluate particular AI against defined benchmark tasks that capture well-understood and clearly defined everyday regulatory activities. While that would give rise to the private sector optimising against those benchmark tasks, constructive ambiguity in their design can hamper such efforts. We have seen many examples in the pre-AI area, such as capital regulations before the global crisis in 2008, where the reported capital of institutions that ultimately failed did not indicate problems. More recently, banking authorities have moved against internal models and towards simple buffers. The authorities need to pay particular attention to how AI facilitate optimisation against models.
Bengio et al. (2023) and Shevlane et al. (2023) in their analysis of the use of AI, propose avenues for benchmarking, such as checking for engine consistency, that is, whether standard inputs yield reasonable outputs, and if similar inputs yield equally similar outputs. The authorities could then execute regular scans for engines with unexpected capabilities and test them to see if they can engage in manipulative behaviour and if they are transparent, via explainability, in how they reach conclusions. Ultimately, transparency, for example, by publishing sufficient details on architectures and training or even by using open-source approaches, will allow outside scrutiny. Whether that is sufficient remains an open question.
5.4 Regulation of public AI
The financial authorities act as agents for society and must be held accountable to it. We have several safeguards built to achieve that, including governance structures and judicial and parliamentary scrutiny. As the authorities increasingly start using AI, new areas of accountability emerge.
We have several ways today to provide accountability. The human supervisor who makes mistakes can be held responsible, re-trained, dismissed or even prosecuted, as we have seen many times in the past. If AI is tasked with making the same type of supervisory decisions, how does one attribute accountability? While the supervisory structure has many individuals, where perhaps only one makes a mistake, the authority might only use one AI engine to make all supervisory decisions. To overcome that, it may be necessary to run multiple engines, but that creates its own problems when their analyses conflict.
When a human supervisor makes a mistake, or the regulated thinks a human supervisor makes a mistake, the regulated can appeal and challenge the ruling in a court of law. That means a human supervisor will be called on to explain the logic behind a decision, and the authority will explain the guidelines given to the human supervisor. That will be more difficult when AI makes decisions. There will be no separation between individual decisions and the guidelines, and the regulatory AI may not be able to explain its reasoning or why it thinks it complies with laws and regulations.
A useful way to evaluate AI use in the public sector would be to test it for safety and performance on defined benchmark tasks and tabletop exercises, particularly activities the authority engages in. It could be asked to advise and rule on macroprudential actions and suggest how to respond to different hypothetical crisis scenarios.
6 Criteria for AI use in financial policy
@@
The issues we discuss above take us to six criteria for evaluating AI use in financial policy.
@@
Can AI be given clear objectives and its actions monitored in light of those objectives, or are they unclear?
@@
Would a human functionary have the authority to make decisions, does it require committee approval, or is a fully distributed decision-making process brought to bear on a problem?
@@
Task Data Mutability Objectives Authority Responsibility Consequences
Fraud/Compliance Consumer protection Ample Very low Clear Single Mostly clear Small
Microprudential risk management Routine forecasting Ample Very low Mostly clear Single Clear Moderate
Criminality Terrorism Limited Very low Mostly clear Multiple Moderate Moderate
Nation state atttacks Limited Full Complex Multiple & international Moderate Very severe
Resolution of small bank failure Limited Partial Clear Mostly single Mostly clear Moderate
Resolution of large bank failure Severe market turmoil Rare Full Complex Multiple Often unclear Severe
Global systemic crises Very rare or not available Full Complex & conflicting Multiple & international Unclear even ex-post Very severe
@@ 7 Conclusion
The literature on AI has identified several areas where it can threaten society, and we augment those with issues arising in the financial system. We find four channels: malicious use, misinformed use and overreliance, misalignment and market structure.
@@ Appendix A Machine Learning (ML)
There are many excellent sources on the technical details of ML, see for example Hastie et al. (2009) and Murphy (2023). The ML algorithms that underpin the current generation of successful AI engines are based on neural networks and involve inferring a complex mapping from some input data to an output. [^13] The algorithms look for a function $f$ that “best” maps input $x$ to output $y$, where $f$ involve multiple nested layers, transforming the input data to extracting complex patterns. The techniques used to fit or train these types of models are referred to as deep learning. See Prince (2023) for an up-to-date overview of deep learning techniques. For example, for most large language models (LLMs), the input is a sequence of what is called tokens, perhaps words, images or sounds, while the output is the next token in the sequence, such as a word in a sentence. [^14] Currently, the most successful architecture for foundation models is the transformer (Vaswani et al., 2023) used in OpenAI’s GPT models, Google’s PaLM and Meta’s LLaMA. While the transformer architecture dominates the current AI landscape, some alternative architectures for foundation models have recently emerged, e.g., state space sequence models (SSMs) (Gu et al., 2022) that are computational efficient and can model long-term dependencies in time-series data.